8/Dec/99 Day Shift

Herman B. White (hwhite@fnal.gov)
Wed, 8 Dec 1999 15:23:01 -0600 (CST)


Shift Summary for eve shift, 7/Dec/99

On Shift: H. Nguyen, H. White

Hours of beam available at all intensities: 0.00
Hours of beam available at nominal intensity: 0.00
Hours of E832 data (special runs not included): 0.00
Hours of special runs (laser scans, muon runs...): 8.00
========================================================================
Accounting of lost time: Magnet change

========================================================================
Runs ended this shift:
Runs 15157-15158 not recorded in the run log.
Current cosmic run 15159

Tapes brought to FCC this shift: None

=======================================================================
Accesses during shift: Open access.
Removed DCOR latch cables to deep latches in crate 51.
Techs replace ion gauge
Tech replaces oil in rotary pumps
Surveyors work on "level run" throughout the day
R. Coleman, R. Ford, G. Tassoto, check out new accidental counter
readout
========================================================================
Overview of shift: No beam during replacement of magnet at D4.

========================================================================
Description of Shift:


run spill time comment
------------------------------------------------------------------------
15159 07:00 Start COSMIC run
DCOR latch cables removed entirely from deep
latches in Crate 51

08:00 Survey crew begin level run
10:00 Tech replace ion gauge
10:15 JAWS restored DC voltage to CsI chan# 2820
had +31.3% gain drift wrt DB voltage
10:29 Rotary oil replacement
12:20 Walk through. No activity in hall
12:40 ktevana errors
13:26 Oil exchange complete
14:35 New accidental counter installed and checked out
(R. Coleman, R. Ford, G. Tassoto)

========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift